OpenStudy (loser66):

Please help prove 4^n >n^2 for all non-negative integers n

OpenStudy (loser66):

Ignore basic step and hypothesis one, I got stuck at the last step only \[4^{n+1}>(n+1)^2\]

OpenStudy (loser66):

I have class now. Please, leave your guidance here. Thanks in advance ;)

OpenStudy (phi):

I was thinking take the log of both sides n log4 > 2 log n (1/2) log4 > (1/n) log n log 2 > log( n^(1/n) ) 2 > n^(1/n) we know 4^n > n^2 for n=0, n=1... the limit as n->infinity of n^(1/n) is 1

OpenStudy (phi):

inductive reasoning: base cases: true for n=0,1,2 assume true for n≥2 4^n > n^2 then 4 * 4^n > 4n^2 4^(n+1) > 4 n^2 aside: use induction to show 4 n^2 > (n+1)^2 for n>=2 base case: 4*2^2 = 16 > 3^2 16>9 true for n+1: 4 (n+1)^2 > ? ( (n+1) +1)^2 4 n^2 + 8n + 4 > ? n^2 +4n + 4 3 n^2 + 4n > ? 0 true for all n≥2, therefore 4 n^2 > (n+1)^2 for n>=2 we use this result to write 4^(n+1) > 4 n^2 > (n+1)^2 4^(n+1) > (n+1)^2 i.e. true for the n+1 instance, given it's true for the nth instance.

OpenStudy (loser66):

Can I go on this way? Basic step: P(1) := \(4^1> 1^2 \\4>1\) P(1) is true Inductive steps: Assume P(n) is true P(n) := \(4^n > n^2\) We need to prove P(n) -> P(n+1) P(n+1) := \(4^{n+1} > (n+1)^2\) the left hand side: \(4^{n+1}= 4^n*4=4^n+4^n+4^n+4^n\) the right hand side: \((n+1)^2 = n^2+2n+1 = n^2+n+n+1\) I have: \(4^n > n^2~~~~~~~~~ \text {(hypothesis step)}\) \(4^n >n ~~~ \forall n>0 \) \(\color{red}{\text {do I have to prove it??}}\) \(4^n>n\) \(4^n >1 \) -------------------------------------------------------------- \(4*4^n > n^2 +2n +1\\4^{n+1}>(n+1)^2\)

OpenStudy (phi):

you could just make a short note: given n>1, multiply both sides by n: n^2 > n therefore 4^n > n^2 implies 4^n > n for n>1

OpenStudy (loser66):

Thank you, I got it. :)